# Mil2: Efficient Cloth Simulation Using Non-distance Barriers and Subspace Reuse

### v3

@@ 9. Conclusion & Limitation
Mil2 is a parallelizable cloth simulation framework, that delivers high-quality animation results, keeps computations lightweight, and separates all the triangles on the cloth models. To the best of our knowledge, such combined efficiency, quality, and performance are not possible with existing cloth animation algorithms. To achieve this goal, Mil2 employs a non-distance barrier formulation, which is both simple and effective. This new barrier model allows the simulation to skip the computation of the distance between primitive pairs, and the stiffness of each collision constraint becomes self-adjusting during LG iterations given the current active set. The subspace reuse scheme significantly pushes the performance of the solver with minimum costs. By observing the fact that the low-frequency subspace is less sensitive to high-frequency collisional deformations, we reuse the rest-shape modal global matrix to solve global cloth deformation. The subspace matrix update is also efficient and can be pre-computed. The residual forwarding helps mitigate the dumpling artifacts due to small-TOI line searches. We show that Mil2 yields visually plausible animations with the penetration-free guarantee and makes the simulation interactive for high-resolution scenes.
@@
Mil2 is not perfect. It also has some limitations, leaving us interesting future research opportunities. First, Mil2 is based on projective dynamics. That said, the elasticity model ($\Psi$) should be shaped in a quadratic form. This prevents us from incorporating more complex fabric models in the simulation e.g., homogenized (Sperl et al., 2020) or data-driven models (Feng et al., 2022). However, we believe it is possible to combine our subspace reuse strategy with block descent methods (Lan et al., 2023) for general materials. The residual forwarding trick used in Mil2 becomes less effective if the residual errors are too large. In this situation, RF can introduce unnatural bumpy artifacts, which may be even more visually disturbing. Nevertheless, we do believe Mil2 echoes a general philosophy for developing post-Moore simulators as an integration of mutually complementary numerical procedures crafted for a specific problem. This enables deeper optimization of the simulation and stronger synergy with emerging computing hardware.

### v4

@@ 9. Conclusion & Limitation
This paper presents a parallelizable cloth simulation framework, that delivers high-quality animation results, keeps computations lightweight, and separates all the triangles on the cloth models. To the best of our knowledge, such combined efficiency, quality, and performance are not possible with existing cloth animation algorithms. To achieve this goal, we employ a non-distance barrier, which is both simple and effective. This new barrier model allows the simulation to skip the computation of the distance between primitive pairs, and the weight of each collision constraint becomes self-adjusting during LG iterations given the current active set. The subspace reuse scheme significantly pushes the performance of the solver with minimum costs. By observing the fact that the low-frequency subspace is less sensitive to high-frequency collisional deformations, we reuse the rest-shape modal global matrix to solve global cloth deformation. The subspace matrix update is also efficient and can be pre-computed. The residual forwarding helps mitigate the dumpling artifacts due to small-TOI line searches. We show that this approach yields visually plausible animations with the penetration-free guarantee and makes the simulation efficient for high-resolution scenes.
Table 2. More statistics on parameters. We report total LG iterations and time cost (in ms) using different simulation parameters. Detailed results (including artifacts and simulation failure cases) are presented in the supplementary videos.
Cloth folding
${\Delta t}(\text{s})$ $\|\Delta x\|=1\cdot 10^{-4}$ $\|\Delta x\|=5\cdot 10^{-4}$ $\|\Delta x\|=1\cdot 10^{-3}$ $\|\Delta x\|=1\cdot 10^{-2}$
$1/150$ $39(16.3\text{ms})$ $18(8.3\text{ms})$ $15(6.9\text{ms})\;$ Jittery $4(2.8\text{ms})$ Jittery
$1/100$ $55(24.7\text{ms})$ $26(13.5\text{ms})$ $20(11.1\text{ms})$ Jittery Fail
$1/50$ $95(56.4\text{ms})$ $45(26.6\text{ms})$ $40(20.9\text{ms})$ Jittery Fail
Cloth twisting
${\Delta t}(\text{s})$ $\|\Delta x\|=1\cdot 10^{-4}$ $\|\Delta x\|=5\cdot 10^{-4}$ $\|\Delta x\|=1\cdot 10^{-3}$ $\|\Delta x\|=1\cdot 10^{-2}$
$1/150$ $214(64.8\text{ms})$ $171(52.2\text{ms})$ $157(47.4\text{ms})$ $72(21.6\text{ms})$
$1/100$ $229(69.7\text{ms})$ $182(56.1\text{ms})$ $163(49.4\text{ms})$ Fail
$1/50$ $255(77.3\text{ms})$ $202(61.1\text{ms})$ $171(51.9\text{ms})$ Fail
Figure 20. Folding and twisting.  As a non-Newton method, the performance of our method is correlated with time step size and termination criteria. RF may introduce unnatural jittery artifacts and may cause simulation failure when the residual error is too large. The animated results are available on supplementary video.
@@
Our method has some limitations. First, our method is based on projective dynamics. Therefore, the elasticity model ($\Psi$) should be shaped in a quadratic form. This prevents us from incorporating more complex fabric models in the simulation e.g., homogenized (Sperl et al., 2020) or data-driven models (Feng et al., 2022). However, we believe it is possible to combine our subspace reuse strategy with block descent methods (Lan et al., 2023) for general materials. Residual forwarding trick becomes less effective if the residual errors are too large and may cause simulation failure: in some situations, RF can introduce unnatural bumpy artifacts, which may be even more visually disturbing (see Tab. 2, two columns on the right, Fig. 20 and supplementary videos). While our method adapts to various simulation parameters, it is more sensitive to timestep size and termination criteria compared to the projection-Newton.
Acknowledgements.
We thank reviewers for their detailed and constructive comments. Yin Yang is partially supported by NSF under grant numbers 2301040, 2008915, 2244651, 2008564. Chenfanfu Jiang is supported in part by NSF CAREER 2153851, CCF 2153863, ECCS-2023780.